xx IDScalarWaveC

IDScalarWaveC

Abstract

1 Introduction 2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters

11amplitude10Scope: privatex01  REAL

11Description: The amplitude of the waves

11 Range 10 x01 Default: 1.0
11*:*  10No restriction

11kx  10Scope: privatex01  REAL

11Description: The wave number in the x-direction

11 Range 10 x01 Default: 4.0
11*:*  10No restriction

11

11ky  10Scope: privatex01  REAL

11Description: The wave number in the y-direction

11 Range 10 x01 Default: 0.0
11*:*  10No restriction

11

11kz  10Scope: privatex01  REAL

11Description: The wave number in the z-direction

11 Range 10 x01 Default: 0.0
11*:*  10No restriction

11

11radius 10Scope: privatex01  REAL

11Description: The radius of the gaussian wave

11 Range 10 x01 Default: 0.0
110:*  10Positive

11

11sigma 10Scope: privatex01  REAL

11Description: The sigma for the gaussian wave

11 Range 10 x01 Default: 0.1
110:*  10Positive

11

11initialxdata 10Scope: restrictedx01  KEYWORD

11Description: Type of initial data

11 Range 10 x01 Default: gaussian
11plane  10Plane wave
11gaussian 10Gaussian wave
11box  10Box wave
11none  10No initial data, zero phi

11

11type  10Scope: shared from GRIDx01KEYWORD

11

7 Interfaces

General Implements: idscalarwave Inherits: wavetoy grid

8 Schedule This section lists all the variables which are assigned storage by thorn CactusWave/IDScalarWaveC. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxPARAMCHECK x idscalarwavecxcheckparameters x check parameters x

11x11Language:11c
11 x 11Type: 11function

11

CCTKxINITIAL x idscalarwavecxinitialdata x initial data for 3d wave equation x

11x11Language:11c
11 x 11Reads: 11grid::x
11 x 11x  11grid::y
11 x 11x  11grid::z
11 x 11Storage: 11wavetoy::scalarevolve[3]
11 x 11Type: 11function
11 x 11Writes: 11wavetoy::phi(everywhere)
11 x 11x  11wavetoy::phixp(everywhere)

Aliased Functions x

11Alias Name: xxxxxxx 11Function Name:
11IDScalarWaveCxInitialData11WaveToyxInitialData